# Towards Modality Generalization: A Benchmark and Prospective Analysis

### v2

@@ Abstract
1 Introduction
@@
The world is perceived and interpreted through diverse modalities, e.g., images, videos, text, and audio [1, 2, 3, 4, 5, 6, 7]. Each modality offers a unique perspective, which captures patterns and describes objects from certain physical viewpoints. Against this backdrop, multi-modal learning has emerged as a powerful approach to developing models capable of processing and integrating information across multiple modalities [8, 9, 10, 11, 12]. Compared to traditional single-modal methods, multi-modal learning not only achieves superior performance in conventional tasks (e.g., object recognition [13] and detection [14]), but also opens up opportunities for exploring novel tasks enabled by cross-modal interactions (e.g., cross-modal retrieval [15], audio-visual learning [16], and visual question answering [17]).
@@
Figure 1: The problem of Modality Generalization (MG). The learner trained on multiple modalities (e.g., video or audio) is capable of performing well on unseen modalities (e.g., depth) during testing.
Aware of the mentioned situation and issue, we highlight the problem of modality generalization (MG) in this paper, which focuses on enabling models to adapt to and perform well on unseen modalities beyond those available during training (see Figure 1). Besides, we consider two practical cases of the problem: Weak MG and strong MG. Specifically, Weak MG refers to the case where some modalities are unseen during training but can be inferred by existing perceptors linked to training modalities. Such perceptors can be realized by many existing modality binding models that are pre-trained with paired modality data to map various modalities (i.e., training and testing modalities) into a joint embedding space [^1] Note that sharing a joint embedding space across different modalities does not inherently address the MG problem. Further explanations are provided in Section 2.2.. By taking advantage of these perceptors, the feature embeddings of unseen modalities can be directly extracted which enables subsequent adaptation and generalization for solving MG. As a comparison, strong MG describes the case where some modalities are neither seen during training nor perceptible by existing models. For example, a model trained on the text and audio data may encounter a new modality like bio-signals, with no perceptors to map it into the joint embedding space. This requires the model to generalize across modalities without direct alignment or shared representations, demanding a higher level of robustness and adaptability.
2 Preliminaries
2.1 Modality Generalization Problem Formulation
Weak MG. In this case, we have some available perceptors for training and testing modalities, which can map them into the joint embedding space for downstream tasks. We denote the perceptors about training modalities as $\{\Phi^{1},\ldots,\Phi^{K}\}$, where $\Phi_{i}$ corresponds to the $i$-th modality. The perceptor about the unseen testing modality is denoted by $\Phi^{K+1}$. Many existing modality-binding models can realize these perceptors, e.g., ImageBind and LanguageBind, which are trained with paired modality data (e.g., image-centric or language-centric) to exploit the joint embedding space across modalities. Afterward, the embeddings of the training and testing modality data can be obtained by $\mathbf{z}^{k}_{i}=\Phi^{k}(\mathbf{x}^{k}_{i})$ and $\mathbf{z}^{K+1}_{i}=\Phi^{K+1}(\mathbf{x}^{K+1}_{i})$. Note that we do not target the improvement of the perceptors and therefore freeze their parameters for embedding extraction. On top of the above models, we define a learner $h$ trained on the embeddings of the training modality data. This learner is then used to generalize to the unseen testing modality for follow-up tasks.
Strong MG. In this situation, we still access the perceptors of training modalities, i.e., $\{\Phi^{1},\ldots,\Phi^{K}\}$ for embedding extraction. Differently, despite various modalities involved with paired data in modality-bind perceptors pre-training, novel modalities emerge with few or even no paired data, exhibiting a more challenging scenario for modality generalization. In this case, we do not have a perceptor for the testing modality to map it into the joint embedding space of training modalities. Therefore, we have to employ some perceptors solely pre-trained within the testing modality data $\mathcal{M}^{K+1}$ for its embedding extraction and finish follow-up tasks with the trained learner together. Here, embedding spaces differ between training and testing modalities, and we consider that Strong MG approaches have the potential to uncover the hidden invariant features while bypassing the embedding space variants. The illustrations about these two cases are shown in Figure 2.
Difference from Weak MG. In Strong MG, the perceptor for the testing modality is trained via self-supervised learning, which provides information relevant to the unseen modality. Therefore, the challenge lies in how the algorithm captures invariances between the training modalities, thus generalizing to the unseen modality. This is more difficult than Weak MG, which builds upon existing modality binding methods.
Potentials of Strong MG. Note that Strong MG is very practical in real-world scenarios where modalities emerge as technology advances. This further inspires us to incorporate Strong MG into our benchmark. In fact, previous works such as [18, 19] also support the potential of Strong MG, where independently trained perceptors of increasing scale can be implicitly aligned with each other.
Figure 2: The framework of ModalBed, involving two cases of Weak MG and Strong MG. These two cases share a common training process, where perceptors $\{\Phi^{1},\dots,\Phi^{K}\}$ extract training modalities $\{\mathcal{M}^{1},\dots,\mathcal{M}^{K}\}$, and a learner $h$ is trained for follow-up tasks. Notably, in weak MG, the testing perceptor $\Phi^{K+1}$ is pre-trained with the same objective as the training perceptors to generate joint embeddings. As a comparison, in strong MG, the testing perceptor is kept isolated from training perceptors.
2.2 Limitations of Joint Embedding Spaces in Addressing the MG Problem
It should be noted that sharing a joint embedding space across different modalities does not inherently solve the MG problem, as it often fails to effectively learn invariant features across modalities for generalization. Invariant features, which capture shared semantic or structural information across different modalities, are essential for generalization. However, it is hard to learn only with joint embedding spaces. For example, if the alignment model lacks sufficient capacity or if the paired data between modalities is sparse, the alignment may focus on local correlations specific to individual modalities rather than extracting globally shared features. Additionally, the inherent heterogeneity between modalities poses significant challenges. For example, the feature space of language data, which emphasizes semantic relationships, differs greatly from that of EEG signals, which encode neural activity patterns. Such heterogeneity makes it difficult for the alignment model to effectively map these modalities into a joint embedding space while preserving their shared invariant features. Without successfully learning these invariant features during training, the shared embedding space cannot generalize well to unseen modalities during testing. This limitation underscores the core challenge of the MG problem, highlighting the need for strategies that explicitly address the learning and transfer of invariant features across diverse and heterogeneous modalities.
2.3 Related Work
Modality binding. The modality bind strategy aims to learn a joint embedding space across different modalities. CLIP [13] pioneers the alignment of image and language data and makes them share an embedding space for downstream tasks, e.g., zero-shot classification [20] and open-vocabulary segmentation [21], and inspires lots of subsequent works on modality alignment [22, 23, 24, 25, 26, 27, 28, 29]. Recently, ImageBind [30] proposes to utilize vision modality as a bridge to unify various modalities into a shared embedding space. LanguageBind [31] employs language as an alternative solution to bridge different modalities similarly. Moreover, UniBind [32] makes the alignment centers modality agnostic and further learns a unified and balanced embedding space, empowered by large language models. Note that modality binding can provide perceptors to some modalities and benefits in solving the MG problem.
Multi-modal learning and generalization. Previous works have focused on the objective of leveraging the knowledge from some modalities and generalizing it to another one. For instance, cross-modal fine-tuning [33, 34, 35] operates similarly to transfer learning by aligning the training modality data distribution with the testing modality data, utilizing a shared model architecture. Cross-modal generalization [36, 37] employs distinct encoders and aims to generalize across different modality data derived from the same instance. Besides, multi-modal learning (MML) w/o labeled multi-modal data [38] proposes a setting with unpaired labeled training modality data and testing modality data, supplemented by unlabeled paired multi-modal data for learning modality interactions. Modality connection [39, 40, 41] aims to solve the mismatched modality across datasets, but it does not explore unseen modalities. Recent out-of-modal generalization studies how to adapt to an unknown modality without given instance-level modal correspondence or scarcely paired correspondence in some cases. However, the setting of our modality generalization is different from theirs, as they require incorporating previously unseen modality data into the training process, whereas our setting does not rely on training with the unseen modality. To further distinguish the problem settings between ours and existing works, we denote the modal knowledge possessed by pre-trained perceptors as “known”, and the modalities shown in the training set as “seen”, otherwise denoted as “unknown” and “unseen”, respectively, as shown in Table 1.
Domain generalization. The process of training a model on multiple source domains to enable generalization to unseen target domains is known as domain generalization [42]. Compared to domain adaption [43], domain generalization is more challenging because target domain data cannot be accessed during training. Previous studies have handled domain generalization from a series of perspectives [42, 35], e.g., data manipulation [44, 45, 46], learning domain-invariant representations [47, 48, 49], using gradient operation [50], with meta-learning [51], and etc. Readers can refer [42] for more details on domain generalization. It is apparent that domain generalization and modality generalization share similar objectives in enabling models to generalize to unseen settings. Therefore, many methodologies developed for domain generalization, such as invariant learning and adversarial alignment, can inspire solutions for modality generalization. In addition, the benchmark in domain generalization (e.g., DomainBed [52]) can help the implementation of modality generalization.
@@ 2.3 Related Work
Table 1: A comparison of different multi-modal learning and generalization problems. Here “Correspondence” indicates instance-level modal correspondence, i.e., multi-modal data are paired with the same instance. We use “✓” to denote the correspondence is needed or employ “✗” otherwise.
@@
Cross-Modal Fine-Tuning [33, 34, 35] Known Known & Seen
Cross-Modal Generalization [36, 37] Known Unknown & Seen
MML w/o labeled Multi-Modal Data [38] Known Unknown & Seen
Modality Connection [39, 40, 41] Known Known & Seen
Out-of-Modal Generalization [53] Known Unknown & Seen
@@
3 ModalBed: A PyTorch Tester for Modality Generalization
Continuing ModalBed. ModalBed is believed to be a valuable tool for benchmarking modality generalization algorithms. Its continued development is essential for advancing the research field. Future updates to ModalBed include: (1) integration of new datasets and modalities: regularly update the platform with datasets involving novel modalities or their combinations to reflect realistic challenges; (2) implementation of cutting-edge algorithms: expand the repository to include emerging techniques in multi-modal and domain/modality generalization research; (3) collaborative contributions: foster an open-source community to encourage researchers to contribute algorithms, datasets, and benchmarks, ensuring ModalBed remains comprehensive and up-to-date.
4 Experiments
4.1 Experimental Setups
@@
Table 2: The important statistics of used experimental datasets. We use “✓” to denote a modality is observable in a dataset or employ “-” otherwise.
@@
Datasets. We consider datasets with at least three modalities: (1) MSR-VTT [54], which includes videos and text descriptions. Here we break down the videos into video frames and the audio data following [55, 32]. (2) NYUDv2 [56], which includes RGB images, depth data, and class names that can be transformed into language descriptions as well. (3) VGGSound [57], which contains video data, corresponding sound, and language descriptions. Note that we randomly sample instances from VGGSound to construct a smaller-scale dataset, enabling more efficient benchmarking. We name this smaller-scale set as VGGSound-S. The important statistics of these used datasets are shown in Table 2. More datasets will be included in future updates.
@@
Perceptors. We utilize three types of perceptors to process different modalities: (1) ImageBind [30]; (2) LanguageBind [31]; (3) UniBind [32]. Generally speaking, with paired data, these perceptors are obtained by learning a unified representation space across different modalities. Specifically, ImageBind leverages the binding property of images and learns a unified shared representation space by utilizing multiple types of image-paired data. By aligning each modality’s embedding to image embeddings, ImageBind achieves emergent alignment across diverse modalities without requiring all modalities to co-occur. Afterward, LanguageBind takes the language as the bind across different modalities and aligns the embeddings of the other modalities to corresponding language embeddings, implementing multi-modal semantic alignment. Furthermore, UniBind does not treat the image or language as the central modality. Differently, it makes the alignment centers modality agnostic and further learns a unified and balanced representation space. For weak modality generalization, the perceptors for both training and testing are available and sourced from the modality bind method. For strong modality generalization, the perceptors for training modalities still come from the modality bind. The perceptor of the testing modality is obtained by self-supervised learning only with the data of this modality.
In this work, we use contrastive learning [58] to implement self-supervised learning. Specifically, we mask parts of the loaded data. Text data is represented as token indices, and other modalities are loaded as matrices, with a masking ratio of 0.3 to generate two views for each instance. Masked elements are set to zero. We then apply contrastive learning to pull these two views closer in the representation level.

### v3

@@ Abstract.
Multimodal Learning, Modality Generalization
copyright: acmlicensedjournalyear: 2025copyright: ccconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia (MM ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3758174isbn: 979-8-4007-2035-2/2025/10ccs: Computing methodologies Computer vision representationsccs: Theory of computation Machine learning theory
1. Introduction
@@
The world is perceived and interpreted through diverse modalities, e.g., images, videos, text, and audio (Wu et al., 2024b; Zhang et al., 2023a; Lu, 2024, 2023; Wu et al., 2024a; Luo et al., 2024b, a; Zhou et al., 2025a, 2024b, b). Each modality offers a unique perspective, which captures patterns and describes objects from certain physical viewpoints. Against this backdrop, multi-modal learning has emerged as a powerful approach to developing models capable of processing and integrating information across multiple modalities (Huang et al., 2021; Xu et al., 2023; Ektefaie et al., 2023; Bayoudh et al., 2022; Liang et al., 2024; Luo et al., 2025). Compared to traditional single-modal methods, multi-modal learning not only achieves superior performance in conventional tasks (e.g., object recognition (Radford et al., 2021) and detection (Chen et al., 2022)), but also opens up opportunities for exploring novel tasks enabled by cross-modal interactions (e.g., cross-modal retrieval (Zhen et al., 2019), audio-visual learning (Zhu et al., 2021), and visual question answering (Cadene et al., 2019)).
@@
Figure 1. The problem of Modality Generalization (MG). The learner trained on multiple modalities (e.g., video or audio) is capable of performing well on unseen modalities (e.g., depth) during testing.
@@
Table 1. A comparison of different multi-modal learning and generalization problems. Here “Correspondence” indicates instance-level modal correspondence, i.e., multi-modal data are paired with the same instance. We use “✓” to denote the correspondence is needed or employ “✗” otherwise.
@@
Cross-Modal Fine-Tuning (Cai et al., 2024; Shen et al., 2023; Dong et al., 2023) Known Known & Seen
Cross-Modal Generalization (Liang et al., 2021; Xia et al., 2024) Known Unknown & Seen
MML w/o labeled Multi-Modal Data (Liang et al., 2023) Known Unknown & Seen
Modality Connection (Wang et al., 2023b, a; Ma et al., 2019) Known Known & Seen
Out-of-Modal Generalization (Huang et al., 2025) Known Unknown & Seen
@@
Aware of the mentioned situation and issue, we highlight the problem of modality generalization (MG) in this paper, which focuses on enabling models to adapt to and perform well on unseen modalities beyond those available during training (see Figure 1). Besides, we consider two practical cases of the problem: Weak MG and strong MG. Specifically, Weak MG refers to the case where some modalities are unseen during training but can be inferred by existing perceptors linked to training modalities. Such perceptors can be realized by many existing modality binding models that are pre-trained with paired modality data to map various modalities (i.e., training and testing modalities) into a joint embedding space [^1] Note that sharing a joint embedding space across different modalities does not inherently address the MG problem. Further explanations are provided in Section 3.4.. By taking advantage of these perceptors, the feature embeddings of unseen modalities can be directly extracted which enables subsequent adaptation and generalization for solving MG. As a comparison, strong MG describes the case where some modalities are neither seen during training nor perceptible by existing models. For example, a model trained on the text and audio data may encounter a new modality like bio-signals, with no perceptors to map it into the joint embedding space. This requires the model to generalize across modalities without direct alignment or shared representations, demanding a higher level of robustness.
Figure 2. The framework of ModalBed, involving two cases of Weak MG and Strong MG. These two cases share a common training process, where perceptors $\{\Phi^{1},\dots,\Phi^{K}\}$ extract training modalities $\{\mathcal{M}^{1},\dots,\mathcal{M}^{K}\}$, and a learner $h$ is trained for follow-up tasks. Notably, in Weak MG, the testing perceptor $\Phi^{K+1}$ is pre-trained with the same objective as the training perceptors to generate joint embeddings. As a comparison, in Strong MG, the testing perceptor is kept isolated from training perceptors.
2. Related Work
2.1. Modality Binding
The modality binding strategy aims to learn a joint embedding space across different modalities. CLIP (Radford et al., 2021) pioneers the alignment of image and language data and makes them share an embedding space for downstream tasks, e.g., zero-shot classification (Qian and Hu, 2024) and open-vocabulary segmentation (Wang et al., 2024), and inspires lots of subsequent works on modality alignment (Li et al., 2022, 2023; Song et al., 2023; Guo et al., 2023; Yang et al., 2024a; Zhou et al., 2024a; Zhang et al., 2023b; Liu et al., 2025a, b). Recently, ImageBind (Girdhar et al., 2023) proposes to utilize vision modality as a bridge to unify various modalities into a shared embedding space. LanguageBind (Zhu et al., 2024) employs language as an alternative solution to bridge different modalities similarly. Moreover, UniBind (Lyu et al., 2024) makes the alignment centers modality agnostic and further learns a unified and balanced embedding space, empowered by large language models. Note that modality binding can provide perceptors to some modalities and benefits in solving the MG problem.
2.2. Multi-Modal Learning and Generalization
Previous works have focused on the objective of leveraging the knowledge from some modalities and generalizing it to another one. For instance, cross-modal fine-tuning (Cai et al., 2024; Shen et al., 2023; Dong et al., 2023) operates similarly to transfer learning by aligning the training modality data distribution with the testing modality data, utilizing a shared model architecture. Cross-modal generalization (Liang et al., 2021; Xia et al., 2024) employs distinct encoders and aims to generalize across different modality data derived from the same instance. Besides, multi-modal learning (MML) w/o labeled multi-modal data (Liang et al., 2023) proposes a setting with unpaired labeled training modality data and testing modality data, supplemented by unlabeled paired multi-modal data for learning modality interactions. Modality connection (Wang et al., 2023b, a; Ma et al., 2019) aims to solve the mismatched modality across datasets, but it does not explore unseen modalities. Recent out-of-modal generalization studies how to adapt to an unknown modality without given instance-level modal correspondence or scarcely paired correspondence in some cases. However, the setting of our modality generalization is different from theirs, as they require incorporating previously unseen modality data into the training process, whereas our setting does not rely on training with the unseen modality. To further distinguish the problem settings between ours and existing works, we denote the modal knowledge possessed by pre-trained perceptors as “known”, and the modalities shown in the training set as “seen”, otherwise denoted as “unknown” and “unseen”, respectively, in Table 1.
2.3. Domain Generalization
The process of training a model on multiple source domains to enable generalization to unseen target domains is known as domain generalization (Wang et al., 2022). Compared to domain adaption (Pan et al., 2010), domain generalization is more challenging because target domain data cannot be accessed during training. Previous studies have handled domain generalization from a series of perspectives (Wang et al., 2022; Dong et al., 2023), e.g., data manipulation (Tobin et al., 2017; Zhang, 2017; Zhou et al., 2020), learning domain-invariant representations (Ganin et al., 2016; Li et al., 2018a; Lin et al., 2022), using gradient operation (Huang et al., 2020), with meta-learning (Li et al., 2018b), and etc. Readers can refer (Wang et al., 2022) for more details on domain generalization. Apparently, domain generalization and modality generalization share similar objectives in enabling models to generalize to unseen settings. Many methodologies developed for domain generalization, such as invariant learning and adversarial alignment, can inspire solutions for modality generalization. In addition, the benchmark in domain generalization (e.g., DomainBed (Gulrajani and Lopez-Paz, 2021)) can help modality generalization implementation.
3. Preliminaries
3.1. Problem Formulation
3.2. Weak MG
In this case, we have some available perceptors for training and testing modalities, which can map them into the joint embedding space for downstream tasks. We denote the perceptors about training modalities as $\{\Phi^{1},\ldots,\Phi^{K}\}$, where $\Phi_{i}$ corresponds to the $i$-th modality. The perceptor about the unseen testing modality is denoted by $\Phi^{K+1}$. Many existing modality-binding models can realize these perceptors, e.g., ImageBind and LanguageBind, which are trained with paired modality data (e.g., image-centric or language-centric) to exploit the joint embedding space across modalities. Afterward, the embeddings of the training and testing modality data can be obtained by $\mathbf{z}^{k}_{i}=\Phi^{k}(\mathbf{x}^{k}_{i})$ and $\mathbf{z}^{K+1}_{i}=\Phi^{K+1}(\mathbf{x}^{K+1}_{i})$. Note that we do not target the improvement of the perceptors and therefore freeze their parameters for embedding extraction. On top of the above models, we define a learner $h$ trained on the embeddings of the training modality data. This learner is then used to generalize to the unseen testing modality for follow-up tasks.
3.3. Strong MG
In this situation, we still access the perceptors of training modalities, i.e., $\{\Phi^{1},\ldots,\Phi^{K}\}$ for embedding extraction. Differently, despite various modalities involved with paired data in modality-bind perceptors pre-training, novel modalities emerge with few or even no paired data, exhibiting a more challenging scenario for modality generalization. In this case, we do not have a perceptor for the testing modality to map it into the joint embedding space of training modalities. Therefore, we have to employ some perceptors solely pre-trained within the testing modality data $\mathcal{M}^{K+1}$ for its embedding extraction and finish follow-up tasks with the trained learner together. Here, embedding spaces differ between training and testing modalities, and we consider that Strong MG approaches have the potential to uncover the hidden invariant features while bypassing the embedding space variants. The illustrations about these two cases are shown in Figure 2.
$\bullet$ Difference from Weak MG. In Strong MG, the perceptor for the testing modality is trained via self-supervised learning, which provides information relevant to the unseen modality. Therefore, the challenge lies in how the algorithm captures invariances between the training modalities, thus generalizing to the unseen modality. This is more difficult than Weak MG, which builds upon existing modality binding methods.
$\bullet$ Potentials of Strong MG. Note that Strong MG is very practical in real-world scenarios where modalities emerge as technology advances. This further inspires us to incorporate Strong MG into our benchmark. In fact, previous works such as (Tjandrasuwita et al., 2025; Huh et al., 2024) also support the potential of Strong MG, where independently trained perceptors of increasing scale can be implicitly aligned with each other.
@@ 3.3. Strong MG
Table 2. The important statistics of experimental datasets. We use “✓” for observable modality in a dataset and “-” otherwise.
@@
3.4. Limitations of Joint Embedding Spaces
Note that sharing a joint embedding space across different modalities does not inherently solve the MG problem, as it often fails to effectively learn invariant features across modalities for generalization. Invariant features, which capture shared semantic or structural information across different modalities, are essential for generalization. However, it is hard to learn only with joint embedding spaces. For example, if the alignment model lacks sufficient capacity or if the paired data between modalities is sparse, the alignment may focus on local correlations specific to individual modalities rather than extracting globally shared features. Besides, the inherent heterogeneity between modalities poses significant challenges. For example, the feature space of language data, which emphasizes semantic relationships, differs greatly from that of EEG signals, which encode neural activity patterns. Such heterogeneity makes it difficult for the alignment model to map these modalities into a joint embedding space while preserving their shared invariant features. Without successfully learning these invariant features during training, the shared embedding space cannot generalize well to unseen modalities during testing. This limitation underscores the core challenge of the MG problem, highlighting the need for strategies that explicitly address the learning and transfer of invariant features across diverse and heterogeneous modalities.
4. ModalBed: A PyTorch Tester for Modality Generalization
$\bullet$ Continuing ModalBed. ModalBed is believed to be a valuable tool for benchmarking modality generalization algorithms. Its continued development is essential for advancing the research field. Future updates to ModalBed include: (1) integration of new datasets and modalities: regularly update the platform with datasets involving novel modalities or their combinations to reflect realistic challenges; (2) implementation of cutting-edge algorithms: expand the repository to include emerging techniques in multi-modal and domain/modality generalization research; (3) collaborative contributions: foster an open-source community to encourage researchers to contribute algorithms, datasets, and benchmarks, ensuring ModalBed remains comprehensive and up-to-date.
5. Experimental Setups
5.1. Datasets
We consider datasets with at least three modalities: (1) MSR-VTT (Xu et al., 2016), which includes videos and text descriptions. Here we break down the videos into video frames and the audio data following (Akbari et al., 2021; Lyu et al., 2024). (2) NYUDv2 (Silberman et al., 2012), which includes RGB images, depth data, and class names that can be transformed into language descriptions as well. (3) VGGSound (Chen et al., 2020b), which contains video data, corresponding sound, and language descriptions. Note that we randomly sample instances from VGGSound to construct a smaller-scale dataset, enabling more efficient benchmarking. We name this smaller-scale set as VGGSound-S. The important statistics of these used datasets are shown in Table 2. More datasets will be included in future updates.
@@ 5.1. Datasets
5.2. Perceptors
We utilize three types of perceptors to process different modalities: (1) ImageBind (Girdhar et al., 2023); (2) LanguageBind (Zhu et al., 2024); (3) UniBind (Lyu et al., 2024). Generally speaking, with paired data, these perceptors are obtained by learning a unified representation space across different modalities. Specifically, ImageBind leverages the binding property of images and learns a unified shared representation space by utilizing multiple types of image-paired data. By aligning each modality’s embedding to image embeddings, ImageBind achieves emergent alignment across diverse modalities without requiring all modalities to co-occur. Afterward, LanguageBind takes the language as the bind across different modalities and aligns the embeddings of the other modalities to corresponding language embeddings, implementing multi-modal semantic alignment. Furthermore, UniBind does not treat the image or language as the central modality. Differently, it makes the alignment centers modality agnostic and further learns a unified and balanced representation space. For weak modality generalization, the perceptors for both training and testing are available and sourced from the modality bind method. For strong modality generalization, the perceptors for training modalities still come from the modality bind. The perceptor of the testing modality is obtained by self-supervised learning only with the data of this modality. In this work, we use contrastive learning (Chen et al., 2020a) to implement self-supervised learning. Specifically, we mask parts of the loaded data. Text data is represented as token indices, and other modalities are loaded as matrices, with a masking ratio of 0.3 to generate two views for each instance. Masked elements are set to zero. We then apply contrastive learning to pull these two views closer in the representation level.
5.3. Algorithms